At a place where the acceleration due to gravity is 10 m/sec^2 a force of 5kg weight acts on a body of mass 10 kg initially at rest. The velocity of the body after 4 sec is?Please give full answer

Grade:9

1 Answers

Nimish
6 years ago
F= ma Can also be written as ,F= mv/tPutting the values,5×10=10×v/4Solving the equation we getv=20m/sec